Seismic activity at the volcano today has again been relatively low despite a slight increase in the number of earthquakes. Between 16:00 yesterday and 16:00 today there were 29 hybrid earthquakes, 15 rockfall signals and 10 long-period earthquakes. No volcano-tectonic earthquakes were recorded.
Most of the seismic activity occurred during two periods; the first extended from 9:00 pm last night to 1:00 am this morning, while the second occurred from 10:00 am to 2:00 pm today.
Visibility has been poor today, and no good views of the dome were obtained. However field teams working on the Boxing Day deposits in the southern part of the island observed several small rockfalls from the active areas of dome growth in the south-western part of English's Crater.
The results of the GPS survey of sites at Harris, Whites, Long Ground, Windy Hill and Brodericks indicated some minor change at the Brodericks station. This is only the second occupation of the Brodericks site and the measurements would have to be repeated to determine whether the changes are real. An EDM survey of the Lees-MVO-Garibaldi Hill triangle was again completed today. This triangle has consistently shown some minor fluctuations in line lengths to the Lees Yard triangle since measurements stated in July 1997. These changes are due to instability at the reflector site and are not considered significant.
Field investigation and sampling of the debris avalanche, pyroclastic flow and surge deposits from the Boxing Day eruption continued today. These results would be used along with previous investigations to help determine the sequence of events and cause of the December 26th collapse. The results of this investigation would be presented as an MVO Special Scientific Report at a later date.
Dust levels at all of the monitoring stations across the island were low today.
People should remain alert and listen to Radio Montserrat for further information.